OpenStudy (anonymous):

c=5/9 (f-32) Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 10 degrees Celsius is.

OpenStudy (anonymous):

Part 1: c = (5/9) (f-32) Multiply both sides of the equation by 9/5 (9/5)c = f - 32 Add 32 to both sides (9/5)c + 32 = f Part 2: (9/5)*10 + 32 = f 18*10 + 32 = f 180 + 32 = f 212 = f